# 3D Shape Completion on Unseen Categories: A Weakly-supervised Approach

### v1

@@ IV-C5 Relationship between the training results and hyperparameters
During the training process, our CSRN determines hyperparameters by evaluating the IoU between the predicted complete shape and the partial input. This approach is able to ensure that the selected hyperparameters are optimized for achieving the highest IoU. Figure 6 illustrates the impact of different hyperparameters on the training results. We discovered that the best training results were obtained when simultaneously using $L_{s}$ and $L_{v}$ and $\lambda=0.5$. Moreover, the rankings of the training results align with the rankings of the test results across all categories, as shown in Table VII. These findings affirm the rationality of selecting optimal hyperparameters based on the IoU between the predicted complete shape and the partial input.
@@ IV-D Ablation Study
IV-D1 Effects of the Prior-assisted Shape Learning Network and Category-specific Shape Refinement Network
To evaluate the effectiveness of the PSLN and CSRN, we conduct ablation studies on both ShapeNet and ScanNet. As Table VI shows, without the Multi-scale Structure Learning unit, the PSLN’s performance falls behind that of PatchComplete on both datasets. Nevertheless, after using the MSL unit, the model improves the IoU by 0.032 and 0.054 in ShapeNet and ScanNet, respectively, surpassing that of PatchComplete. This observation highlights the significant role of our MSL unit in learning complete shapes. This can be attributed to the incorporation of the correlations between multi-scale local patterns in the partial shape and priors. Furthermore, after refinement by our CSRN, our model increases the IoU by 0.033 and 0.047 in ShapeNet and ScanNet, respectively. These results prove that both our PSLN and CSRN can achieve good performance.
@@ IV-D1 Effects of the Prior-assisted Shape Learning Network and Category-specific Shape Refinement Network
Figure 8: Visualization of the coarse shapes outputted by PSLN and the results refined by CSRN. (a) Coarse shapes outputted by PSLN (b) Results after refinement with $\lambda=0.5$. (c) Ground truth with complete shape. Note that this figure only shows the missing parts of the coarse shapes and refined shapes since we only use the missing parts of the coarse shapes as weak supervision. The model tends to reconstruct shapes that are more reasonable and close to the ground truths rather than the coarse shapes.
IV-D2 Effects of the components of the Category-specific Shape Refinement Network
@@
As CSRN plays a critical role in our model, we conducted ablation experiments to verify the impact of each component in CSRN, including $L_{s}$ and $L_{v}$ of voxel-based partial matching loss, the weight $\lambda$ of the coarse shape loss and the priors. As shown in Table VII, without $L_{s}$ and $L_{v}$, the model’s performance is significantly poor, resulting in an average IoU of less than 0.1. Upon adding $L_{v}$, the IoU increases slightly. However, after adding $L_{s}$, the IoU improves nearly three times. That means both $L_{s}$ and $L_{v}$ have an impact on the model’s performance, with $L_{s}$ having a considerably larger impact. In order to visually demonstrate the effect of $L_{s}$ and $L_{v}$, we present some qualitative results before and after incorporating $L_{s}$ and $L_{v}$. As Figure 7 shows, the model tends to occupy most of the voxel grids without $L_{s}$ and $L_{v}$. From our perspective, this behavior can be attributed to the lack of occupancy constraints, leading the model to predict most voxels as foregrounds in order to cover all objects’ shapes and thereby minimize the loss.
@@
In addition, based on the data presented in rows 4, 5, and 7 of Table VII, it is evident that the model yields optimal results when $\lambda$ is 0.5. When the supervision of the coarse shape is excessively strong ($\lambda=1$) or absent ($\lambda=0$), the model’s predictive ability is compromised. To provide a more intuitive analysis of the role of the coarse shape loss, we compare the visual results of both the coarse shapes and the outputs refined with $\lambda=0.5$. As Figure 8 shows, employing the coarse shapes as weak supervision resulted in a more precise prediction. Based on the comprehensive quantification and visualization results, using the coarse shape as weak supervision ($\lambda=0.5$) effectively enhances the reconstruction ability of the model. On the one hand, the coarse shapes can offer effective weak supervision to the model. On the other hand, this weak supervision does not confine the model to reconstruct the same results as the coarse shapes but rather facilitates the model learning to infer shapes that are more accurate and complete.
Besides, from the last two rows of Table VII, we can see that the IoU improves by 0.014 when using our category-specific priors instead of the fixed priors designed by [9]. That means our category-specific priors are better suited for refining data from different categories.

### v2

@@ IV-C5 Relationship between the training results and hyperparameters
During the training process, our CaSR determines hyperparameters by evaluating the IoU between the predicted complete shape and the partial input. This approach can ensure that the selected hyperparameters are optimized to achieve the highest IoU. Figure 6 illustrates the impact of different hyperparameters on the training results. We discovered that the best training results were obtained when simultaneously using $L_{s}$ and $L_{v}$ and $\lambda=0.5$. Moreover, the rankings of the training results align with the rankings of the test results across all categories, as shown in Table VII. These findings affirm the rationality of selecting optimal hyperparameters based on the IoU between the predicted complete shape and the partial input.
@@ IV-D Ablation Study
IV-D1 Effects of CoSL and CaSR
To evaluate the effectiveness of the CoSL and CaSR model, we conducted ablation studies on both ShapeNet and ScanNet. As Table VI shows, without the Multi-scale Structure Learning unit, the CoSL’s performance falls behind that of PatchComplete on both datasets. Nevertheless, after using the MSL unit, the model improves the IoU by 0.032 and 0.054 in ShapeNet and ScanNet, respectively, surpassing that of PatchComplete. This observation highlights the significant role of our MSL unit in learning complete shapes. This can be attributed to the incorporation of the correlations between multi-scale local patterns in the partial shape and priors. Furthermore, after refinement by our CaSR, our model increases the IoU by 0.033 and 0.047 in ShapeNet and ScanNet, respectively. These results prove that both our CoSL and CaSR can achieve good performance.
@@ IV-D1 Effects of CoSL and CaSR
Figure 8: Visualization of the coarse shapes outputted by CoSL and the results refined by CaSR. (a) Coarse shapes outputted by CoSL (b) Results after refinement with $\lambda=0.5$. (c) Ground truth with complete shape. Note that this figure only shows the missing parts of the coarse shapes and refined shapes since we only use the missing parts of the coarse shapes as weak supervision. The model tends to reconstruct shapes that are more reasonable and close to the ground truths rather than the coarse shapes.
[…]
Stage Prior size Prior source Average Bag Lamp Bathtub Bed
CoSL 32 Mean-shift 0.385 0.376 0.341 0.474 0.350
64 Mean-shift 0.395 0.374 0.355 0.484 0.365
112 Mean-shift 0.420 0.391 0.380 0.525 0.383
150 Mean-shift 0.394 0.386 0.345 0.488 0.357
112 Random #1 0.382 0.358 0.341 0.480 0.348
112 Random #2 0.399 0.370 0.368 0.503 0.356
112 Novel #1 0.389 0.366 0.353 0.482 0.356
112 Novel #2 0.392 0.370 0.357 0.484 0.356
CaSR 32 Specific 0.426 0.439 0.358 0.525 0.383
64 Specific 0.437 0.442 0.360 0.525 0.420
112 Specific 0.450 0.450 0.380 0.542 0.426
150 Specific 0.437 0.441 0.358 0.525 0.424
IV-D2 Effects of the components of CaSR
@@
As CaSR plays a critical role in our model, we conducted ablation experiments to verify the impact of each component in CaSR, including $L_{s}$ and $L_{v}$ of voxel-based partial matching loss, the weight $\lambda$ of the coarse shape loss and the priors. As shown in Table VII, without $L_{s}$ and $L_{v}$, the model’s performance is significantly poor, resulting in an average IoU of less than 0.1. Upon adding $L_{v}$, the IoU increases slightly. However, after adding $L_{s}$, the IoU improves nearly three times. That means both $L_{s}$ and $L_{v}$ have an impact on the model’s performance, with $L_{s}$ having a considerably larger impact. In order to visually demonstrate the effect of $L_{s}$ and $L_{v}$, we present some qualitative results before and after incorporating $L_{s}$ and $L_{v}$. As Figure 7 shows, the model tends to occupy most of the voxel grids without $L_{s}$ and $L_{v}$. From our perspective, this behavior can be attributed to the lack of occupancy constraints, leading the model to predict most voxels as foregrounds in order to cover all objects’ shapes and thereby minimize the loss.
TABLE IX: Quantitative results at a resolution of 64 on the ShapeNet dataset. The best results under each metric are highlighted in bold. $\uparrow$” (resp. $\downarrow$”) indicates the higher (resp. lower), the better.
IoU $\uparrow$ PatchComplete [9] 0.562 0.490 0.478 0.610 0.504 0.623 0.517 0.676 0.601
Ours 0.598 0.530 0.511 0.662 0.527 0.649 0.544 0.711 0.650
F1 $\uparrow$ PatchComplete [9] 0.710 0.651 0.638 0.753 0.666 0.751 0.676 0.803 0.745
Ours 0.739 0.686 0.666 0.791 0.685 0.769 0.700 0.827 0.783
CD $\downarrow$ PatchComplete [9] 2.65 2.54 3.47 2.29 3.07 2.93 3.08 2.29 1.51
AnchorFormer [57] 3.53 3.14 5.07 2.85 4.33 3.53 3.83 3.12 2.37
Ours 2.59 2.39 3.20 2.19 3.12 2.94 3.09 2.39 1.36
Figure 9: Visual comparison of the results at a resolution of 64. (a) Partial input (b) PatchComplete. (c) Ours. (d) Ground truth.
@@
In addition, based on the data presented in rows 4, 5, and 7 of Table VII, it is evident that the model yields optimal results when $\lambda$ is 0.5. When the supervision of the coarse shape is excessively strong ($\lambda=1$) or absent ($\lambda=0$), the reconstruction ability of the model is compromised. To provide a more intuitive analysis of the role of the coarse shape loss, we compare the visual results of both the coarse shapes and the outputs refined with $\lambda=0.5$. As Figure 8 shows, employing the coarse shapes as weak supervision resulted in a more precise prediction. Based on the comprehensive quantification and visualization results, using the coarse shape as weak supervision ($\lambda=0.5$) effectively enhances the reconstruction ability of the model. On the one hand, the coarse shapes can offer effective weak supervision to the model. On the other hand, this weak supervision does not confine the model to reconstruct the same results as the coarse shapes but rather facilitates the model learning to infer shapes that are more accurate and complete.
Figure 10: Visual comparison of the results with different noise levels. Noises are highlighted with purple points. (a)$\sim$(b) Noisy input with $\sigma=0.14$ and its output. (c)$\sim$(d) Noisy input with $\sigma=0.15$ and its output. (e)$\sim$(f) Noisy input with $\sigma=0.16$ and its output. (g) Ground truth.
Besides, from the last two rows of Table VII, we can see that the IoU improves by 0.014 when using our category-specific priors instead of the fixed priors designed by [9]. That means our category-specific priors are better suited for refining data from different categories. Furthermore, we investigated the impact of both the size and construction method of the prior bank on the results. First, we varied the size of the prior bank to 32, 64, and 150, respectively, and conducted experiments across both the CoSL and CaSR stages. As detailed in Table VIII, using 112 priors achieves the best results for both stages. Reducing the number of priors leads to performance degradation, while increasing the number to 150 also results in decreased performance. In our opinion, this is because the local structures provided by the priors already cover most shapes with 112 objects. Adding more priors does not provide additional useful structural priors for the model but rather introduces irrelevant information into the learning process, which increases the difficulty of computing prior weights and weakens the role of important priors. In addition, we used different methods to construct the prior bank and explore its impact. Concretely, on the one hand, we randomly sampled two sets of priors from the training set. On the other hand, we selected two groups of priors from four categories in the test set (one from basket, printer, and another from laptop, bench) and tested on the remaining four categories. Compared to the original prior bank, these priors differ significantly from the shape distribution of the training set. Table VIII shows that the random priors result in significantly poorer performance compared to the priors obtained from mean-shift clustering. This performance gap arises because clustering ensures more diverse local shapes among priors. Additionally, from rows 3, 7, and 8 of Table VIII, it can be seen that when the shape distribution of priors differs greatly from that of the training set, the performance of the CoSL’s model significantly decreases. Furthermore, according to the ablation results presented in Table VII, CaSR also demonstrates a performance decrease when using priors (fixed priors) that differ from the distribution of the target shapes. We believe that due to significant differences in local structures, the model learns few shape-related structures between partial observations and priors through cross-attention, making it difficult to accurately reconstruct complete shapes.
TABLE X: Effects of the diversity between the training set and test set. “Diverse split” denotes the group with the most significant shape differences between the training set and test set. “Average” means the average IOU of the control group.
Shape diversity Noise diversity
Diverse split Average Clean $\sigma=0.14$ $\sigma=0.15$ $\sigma=0.16$
0.426 0.429 0.441 0.431 0.410 0.370
IV-D3 Effects of the resolution
Higher resolutions usually preserve finer details. To evaluate our model’s performance on high-resolution data, we conducted experiments using data at a resolution of $64\times 64\times 64$. Table IX presents a comparison of the quantitative results among our approach, PatchComplete, and AnchorFormer. Note that converting the voxel grids of $64\times 64\times 64$ resolution into point clouds yields an average of nearly 24,000 points. For a fair comparison, we set the output point number of AnchorFormer to 16,384 and randomly sampled our results to the same number of points when calculating the CD value. As indicated in Table IX, our method significantly outperforms PatchComplete and AnchorFormer at the $64\times 64\times 64$ resolution. Figure 9 visually illustrates these results, showing that our method excels in preserving input details while generating more complete and reasonable shapes.
IV-D4 Evaluation of training/test set diversity
We also investigated the impact of diversity between the training and test sets, including the shape diversity and noise levels, on the model’s performance. Specifically, we first employed an auto-encoder to learn shape codes for all objects. Subsequently, we utilized the K-means algorithm to cluster the average shape codes of the 26 categories. From these clusters, we selected two groups of data with the most significant shape differences as the training set and the test set, respectively. The training set contains 18 categories, i.e., trash bin, bag, bookshelf, cabinet, chair, dishwasher, display, faucet, file cabinet, guitar, lamp, laptop, microwave, piano, printer, stove, table, washing machine, while the test set contains 8 categories, i.e., basket, bed, pots, bathtub, bench, bowl, keyboard, sofa. To ensure a more objective and fair comparison, we also set up a control group. The control group consists of three pairs of training and test sets obtained through random combinations of categories, including the initial category split in Section IV-A. We compared the IOU of the diverse-shape group with the average IOU of the control group. As shown in Table X, the gap between these two groups is only 0.03. This indicates that our method is highly robust to shape diversity between the training set and the test set.
In addition, considering that real-world partial scans may contain noises, we introduced various levels of Gaussian noise ($\mu=0,\sigma=\{0.14,0.15,0.16\}$) to the partial input of the test set to verify its impact on our method. As shown in Table X and Figure 10, the performance of our method is not significantly affected under conditions of low to moderate noise. Significant performance degradation is observed under extremely severe noise levels. This finding validates the robustness of our method in handling noisy data.